# Analytic theory of dropout regularization

### v1

@@ 3.2 Early-time analysis
A first case in which the equations of motion can be solved exactly is in the early-time regime. In this section, we derive the optimal dropout probability in this limit. We assume that the student parameters are initialized with zero norm, i.e., $Q_{ii}=0$, although the calculation can be readily extended to include a finite initial norm. At short times, the overlaps remain close to a symmetric configuration, where each student node is equally aligned with each of the teacher nodes. In other words,
$w_{i}=R\sum_{n=1}^{M}w^{*}_{n}+\tilde{w}_{i}\,,$ (18)
where, with a slight abuse of notation, $R$ denotes a positive scalar and $\tilde{w}_{i}$ is an additional node-dependent random vector, perpendicular to $w^{*}_{n}$ for all $n=1,\ldots,M$. This motivates the ansatz
@@
$\epsilon_{g}\approx\frac{2}{\pi^{2}}\eta^{2}K(M/3+\sigma^{2})(r^{4}+(K-1)r^{3})\alpha-\frac{2}{\pi^{2}}KM\eta r^{2}\alpha\,.$ (22)
@@
In Fig. 2, we compare this early-time result (continuous line) with the exact result obtained by numerically optimizing the full dynamics with $\alpha=0.1$ (crosses), finding good agreement. As expected, the activation probability decreases with increasing label variance $\sigma^{2}$. For large $\sigma^{2}$, we find $r^{*}\approx 2M/(3\eta\sigma^{2})$.
@@
Figure 3: Early-times phase diagram. Above the lines, obtained varying the number $K$ of the student’s hidden nodes, it is beneficial to use dropout, with the optimal activation probability given in Eq. (23). Below the line, it is optimal to avoid dropout, corresponding to activation probability $r=1$.
@@ 3.3 Intermediate-time analysis: unspecialized phase
By numerically integrating the ODEs (10), we observe that the generalization error exhibits a plateau at intermediate times. This plateau corresponds to the unspecialized phase, previously described in Refs. [16, 17, 18, 19, 24]. In this regime, the student weight vectors $w_{i}$ are equally aligned with all teacher weights. This is in contrast to the specialized phase that emerges at later times, where each student node $w_{i}$ specializes by aligning predominantly with a specific teacher weight $w^{*}_{n}$ or becoming inactive. Note that if the second layer is also trained, more complex specialization patterns can emerge, as the network gains additional flexibility in how it distributes representational roles across the hidden units.
In this section, we analytically investigate the role of dropout during the intermediate phase of training. Despite the complexity of the ODEs, exact results can be obtained in the limit of small learning rate $\eta\to 0$, allowing us to derive expressions for both the optimal dropout probability and the corresponding generalization error. We then compute perturbative corrections around this limit to account for finite learning rates. Finally, we examine how dropout influences the transition to specialization. As we will show, this transition only takes place when the activation probability $r$ exceeds a critical threshold $r_{c}$.
@@ 3.3.1 Small-$\eta$ regime
In this section, we investigate how dropout affects the performance of the network in the unspecialized phase. To do so, we identify fixed-point solutions of the ODEs (43)–(44). To make analytic progress, we first consider the small-$\eta$ limit and neglect terms of order $\eta^{2}$ in the ODEs. Notably, this approximation eliminates the dependence on the label noise variance $\sigma^{2}$. We will later perform a perturbative expansion in $\eta$ to recover this dependence to leading order. By setting $f_{R_{in}}=0$ and $f_{Q_{ik}}=0$ to leading order in $\eta$, we find Eqs. (58) and (59) of Appendix A. Since we are interested in an unspecialized solution, we assume the student weights take the form
@@
for some scalar $R_{0}$ to be determined. In other words, for small learning rate $\eta$, we expect the student weight vectors to lie entirely within the subspace spanned by the teacher vectors $\{w^{*}_{n}\}$, without acquiring an additional component in the orthogonal direction, as in Eq. (18). This corresponds to the ansatz
$Q_{ik}=R_{0}^{2}M\,,\qquad R_{in}=R_{0}\,.$ (27)
By inserting this ansatz into Eqs. (58) and (59), we find that it indeed satisfies the fixed-point equations, yielding
$R_{0}=\frac{1}{\sqrt{\left(1+(K-1)r\right)^{2}-M+\sqrt{M^{2}+(K-1)r\left(1+(K-1)r\right)^{2}\left(2+(K-1)r\right)}}}\,.$ (28)
@@
This expression is shown in Fig. 4 as a function of $K$ and $r$ and is in good agreement with the results of numerically integrating the full ODEs (10) at small but finite $\eta$. We observe that at fixed $r$ and $M$, the generalization error at the plateau shows a minimum at a finite value of $K$. Similarly, at fixed $M=2$ and $K=3$, the error is minimized at an intermediate value of the dropout probability, in this case $r^{*}\approx 0.95$.
Figure 4: Left. Generalization error at the plateau vs. the student’s width $K$ at fixed activation probability $r=0.8$. Right. Generalization error at the plateau vs. the activation probability $r$ at fixed student’s width $K=3$. In both panels, the symbols are obtained by numerical integration of Eq. (10) at different values of the learning rate $\eta$ and with $\sigma=0$. The black line marks the theoretical prediction in the small-$\eta$ regime from Eq. (29). The teacher’s width is fixed to $M=2$.
@@
$\epsilon_{g}^{0}\approx\frac{\pi-3}{6\pi}M+\frac{11+12r(r-2)}{24\pi r^{2}}\frac{M}{K^{2}}+\mathcal{O}\left(\frac{1}{K^{3}}\right)\,,$ (31)
@@
Figure 5: Left: Optimal activation probability $r^{*}_{0}$ in the plateau phase as a function of $K-M$ for different values of $M$. The curves are obtained via numerical minimization of the generalization error in Eq. (29), derived in the small-$\eta$ limit. For large $K$, all curves converge to the universal value $r^{*}_{0}=11/12$. Right: Corresponding value of the shifted generalization error, $\epsilon_{g}-\frac{(\pi-3)M}{6\pi}$, as a function of $K-M$.
@@ 3.3.2 Perturbative corrections to the small-$\eta$ regime
Figure 6: Coefficients $a(K,M)$ (left) and $b(K,M)$ (right) as a function of $K-M$ for different values of $M$.
@@ 3.4 Onset of specialization
In this section, we investigate how the specialization transition is affected by the presence of dropout. Recall that in the absence of dropout ($r=1$) this transition has been characterized in Refs. [18, 25, 26], where it was shown that the symmetric, unspecialized phase is linearly unstable. Small perturbations—arising either from noise or from the initial conditions—grow over time, ultimately driving the system toward a specialized phase in which each student node aligns predominantly with a distinct teacher node or is deactivated. For example, in the case $K>M$, a typical specialized configuration takes the form $w_{i}\approx w^{*}_{i}$ for $i\leq M$ and $w_{i}=0$ for $i>M$. The precise dependence of this specialization transition on the initialization of the online dynamics has been investigated in [27]. Note that the hidden-unit specialization transition has also been studied within the asymptotic Bayes-optimal framework, as discussed, for example, in Refs. [28, 29].
@@ 3.5 Long-time behavior
Figure 8: Generalization error as a function of time $\alpha$ for $K=2$, $M=1$, $\eta=0.1$, and different values of node-activation probability $r=0.7,0.8,0.9,0.96,1$. Each panel represents a different value of label noise $\sigma=0.1$ (left), $\sigma=0.7$ (middle), $\sigma=1$ (right). Initial conditions: $Q_{ij}=0\;\forall(i,j)\neq(1,1)$, $Q_{11}=0.1$, $R_{in}=0\;\forall(i,n)\neq(1,1)$, $R_{11}=0.01$.
As shown in the previous section, the long-time behavior depends strongly on the activation probability $r$. For $r<r_{c}$, the plateau phase remains stable and the network stays trapped in this phase even at long times. Conversely, specialization occurs at long times when $r>r_{c}$. As a result, the optimal activation probability $r^{*}$ depends on both the total training time $\alpha$ and the label noise variance $\sigma^{2}$. This behavior is summarized in Fig. 8.
@@ 4 Conclusions
In this paper, we have studied the effects of dropout regularization on the training dynamics of two-layer neural networks in the teacher-student setting. We have derived exact closed-form equations that describe the training dynamics under online SGD in the high-dimensional limit. By analyzing these equations, we have shown that dropout helps reduce detrimental correlations between the nodes of the network. Furthermore, we have studied the optimal dropout probability and the corresponding generalization error, deriving several analytical results at short, intermediate, and long times. Our findings demonstrate that dropout can significantly mitigate the negative effects of label noise, and that the optimal activation probability decreases as the label noise increases. Finally, we have shown analytically that high dropout probabilities prevent the network from reaching a specialized configuration.

### v2

@@ 3.2 Early-time analysis
A first case in which the equations of motion can be solved exactly is in the early-time regime. In this section, we derive the optimal dropout rate in this limit. We assume that the student parameters are initialized with zero norm, i.e., $Q_{ii}=0$, although the calculation can be readily extended to include a finite initial norm. At short times, the overlaps remain close to a symmetric configuration, where each student node is equally aligned with each of the teacher nodes. In other words,
$w_{i}=R\sum_{n}w^{*}_{n}+\tilde{w}_{i}\,$, (19)
where, with a slight abuse of notation, $R$ denotes a positive scalar and $\tilde{w}_{i}$ is an additional node-dependent random vector, perpendicular to $w^{*}_{n}$ for all $n$. This motivates the ansatz
@@
$\epsilon_{g}\approx\frac{2}{\pi^{2}}\eta^{2}K(M/3+\sigma^{2})((K-1)r^{4}+r^{3})\alpha-\frac{2}{\pi^{2}}KM\eta r^{2}\alpha\,.$ (23)
@@
In Fig. 3, we compare this early-time result (continuous line) with the exact result obtained by numerically optimizing the full dynamics with $\alpha=0.1$ (crosses), finding good agreement. As expected, the activation rate decreases with increasing label variance $\sigma^{2}$. For large $\sigma^{2}$ , i.e., for $\sigma^{2}\gg M\max[1/3,32(K-1)/(9\eta)]$ , we find $r^{*}\approx 2M/(3\eta\sigma^{2})$.
@@
Figure 4: Early-times phase diagram. Above the lines, obtained for different values of the number $K$ of hidden nodes in the student, it is beneficial to use dropout, with the optimal activation probability given in Eq. (24). Below the line, it is optimal to avoid dropout, corresponding to activation probability $r=1$.
@@ 3.3 Intermediate-time analysis: unspecialized phase
By numerically integrating the ODEs (10), we observe that the generalization error exhibits a plateau at intermediate times. This plateau corresponds to the unspecialized—or symmetric—phase, previously described in Refs. [16, 17, 18, 19, 24]. This phase is characterized by a symmetry of the student weight vectors $w_{i}$, that are equally aligned with all teacher weights. This is in contrast to the specialized phase that emerges at later times, where each student node $w_{i}$ specializes by aligning predominantly with a specific teacher weight $w^{*}_{n}$ or becoming inactive. Note that if the second layer is also trained, more complex specialization patterns can emerge, as the network gains additional flexibility in how it distributes representational roles across the hidden units.
In this section, we analytically investigate the role of dropout during the intermediate phase of training. Despite the complexity of the ODEs, exact results can be obtained in the limit of small learning rate $\eta\to 0$, allowing us to derive expressions for both the optimal dropout rate and the corresponding generalization error. We then compute perturbative corrections around this limit to account for finite learning rates. Finally, we examine how dropout influences the transition to specialization. As we will show, this transition only takes place when the activation probability $r$ exceeds a critical threshold $r_{c}$.
@@ 3.3.1 Small-$\eta$ regime
In this section, we investigate how dropout affects the performance of the network in the unspecialized phase. To do so, we identify fixed-point solutions of the ODEs (46)–(47). To make analytic progress, we first consider the small-$\eta$ limit and neglect terms of order $\eta^{2}$ in the ODEs. Notably, this approximation eliminates the dependence on the label noise variance $\sigma^{2}$. We will later perform a perturbative expansion in $\eta$ to recover this dependence to leading order. By setting $f_{R_{in}}=0$ and $f_{Q_{ik}}=0$ to leading order in $\eta$, we find Eqs. (61) and (62). Since we are interested in an unspecialized solution, we assume the student weights take the form
@@
for some scalar $R_{0}$ to be determined. In other words, for small learning rate $\eta$, we expect the student weight vectors to lie entirely within the subspace spanned by the teacher vectors $\{w^{*}_{n}\}$, without acquiring an additional component in the orthogonal direction, as in Eq. (19). In other words, in the small-$\eta$ limit the student does not overfit the noise. This can be verified by checking that the corresponding ansatz
$Q_{ik}=R_{0}^{2}M\,,\qquad R_{in}=R_{0}\,$ (28)
is a fixed point of the ODEs. By inserting this ansatz into Eqs. (61) and (62), we find that it indeed satisfies the fixed-point equations, yielding
$\hskip-8.00003pt\frac{1}{R}_{0}={\sqrt{\left(1+(K-1)r\right)^{2}-M+\sqrt{M^{2}+(K-1)r\left(1+(K-1)r\right)^{2}\left(2+(K-1)r\right)}}}\,.$ (29)
@@
This expression is shown in Fig. 5 as a function of $K$ and $r$ and is in good agreement with the results of numerically integrating the full ODEs (10) at small but finite $\eta$. We observe that at fixed $r$ and $M$, the generalization error at the plateau shows a minimum at a finite value of $K$. Similarly, at fixed $M=2$ and $K=3$, the error is minimized at an intermediate value of the dropout rate, in this case $r^{*}\approx 0.95$.
@@
$\epsilon_{g}^{0}\approx\frac{\pi-3}{6\pi}M+\frac{11+12r(r-2)}{24\pi r^{2}}\frac{M}{K^{2}}+\mathcal{O}(\frac{1}{K^{3}})\,$, (32)
@@
Figure 5: Left. Generalization error at the plateau vs. the student’s width $K$ at fixed activation rate $r=0.8$. Right. Generalization error at the plateau vs. the activation rate $r$ at fixed student’s width $K=3$. In both panels, the symbols are obtained by numerical integration of Eq. (10) at different values of the learning rate $\eta$ and with $\sigma=0$. The black line marks the theoretical prediction in the small-$\eta$ regime from Eq. (30). The teacher’s width is fixed to $M=2$.
@@
Figure 6: Left: Optimal activation probability $r^{*}_{0}$ in the plateau phase as a function of KMK-M for different values of $M$. The curves are obtained via numerical minimization of the generalization error in Eq. (30), derived in the small-$\eta$ limit. For large $K$, all curves converge to the universal value $r^{*}_{0}=11/12$. Right: Corresponding value of the shifted generalization error, $\epsilon_{g}-\frac{(\pi-3)M}{6\pi}$, as a function of KMK-M.
@@ 3.3.2 Perturbative corrections to the small-$\eta$ regime
Figure 7: Coefficients $a(K,M)$ (left) and $b(K,M)$ (right) for the plateau phase as a function of KMK-M for different values of $M$.
@@ 3.4 Onset of specialization
In this section, we investigate how the specialization transition is affected by the presence of dropout. Recall that in the absence of dropout ($r=1$) this transition has been characterized in Refs. [18, 25, 26], where it was shown that the symmetric, unspecialized phase is linearly unstable. Small perturbations—arising either from noise or from the initial conditions—grow over time, ultimately driving the system toward a specialized phase in which each student node aligns predominantly with a distinct teacher node or is deactivated. For example, in the case $K>M$, a typical specialized configuration takes the form $w_{i}\approx w^{*}_{i}$ for $i\leq M$ and $w_{i}=0$ for $i>M$. The precise dependence of this specialization transition on the initialization of the online dynamics has been investigated in [27]. Moreover, changing the scaling of the student readout can affect the rate of convergence toward specialization [28]. Note that the hidden-unit specialization transition has also been studied within the asymptotic Bayes-optimal framework, as discussed, for example, in Refs. [29, 30].
@@ 3.5 Long-time behavior
3.5.1 Perturbative expansion for small $\eta$ ($K=M$)
Figure 9: Coefficients $a(M)$ (left) and $b(M)$ (right) as a function of $M$ for the specialized phase with $K=M$.
To investigate the long-time behavior, we first focus on the case $K=M$ and the small-$\eta$ regime. In the limit $\eta\to 0$, perfect specialization, corresponding to a student perfectly reproducing the teacher with vanishing generalization error, is obtained if $r=1$. In this limit, the specialized order parameters are $Q_{ij}=R_{ij}=\delta_{ij}$.
We compute the first-order corrections in $\eta$. The ansatz in this case is
$Q_{ij}=\delta_{ij}+\eta\delta_{ij}Q_{1}+\eta(1-\delta_{ij})C_{1}\,\quad R_{ij}=\delta_{ij}+\eta\delta_{ij}R_{1}+\eta(1-\delta_{ij})S_{1}\,.$ (43)
One can then perform a perturbative calculation identical to that in Section 3.3.2. We find that in the small-$\eta$ limit the optimal activation probability is given by
$r^{*}\approx\min\left[1;\,1+\eta(a(M)+\sigma^{2}b(M))\right]\,$, (44)
The coefficients $a(M)$ and $b(M)$ are shown in Fig. 9 and can be computed for arbitrary $M$ via the corresponding Mathematica notebook at [23]. Once again, we observe that the coefficient $b(M)$ is negative, implying that larger noise $\sigma$ requires setting lower values of the activation probability.
3.5.2 Numerical analysis
Figure 10: Generalization error as a function of time $\alpha$ for $K=2$, $M=1$, $\eta=0.1$, and different values of node-activation rate $r=0.7,0.8,0.9,0.96,1$. Each panel represents a different value of label noise $\sigma=0.1$ (left), $\sigma=0.7$ (middle), $\sigma=1$ (right).
As shown in section 3.4, the long-time behavior depends strongly on the activation probability $r$. For $r<r_{c}$, the plateau phase remains stable and the network stays trapped in this phase even at long times. Conversely, specialization occurs at long times when $r>r_{c}$. As a result, the optimal activation probability $r^{*}$ depends on both the total training time $\alpha$ and the label noise variance $\sigma^{2}$. This behavior is summarized in Fig. 10.
@@ 4 Conclusions
In this paper, we have studied the effects of dropout regularization on the training dynamics of two-layer neural networks in the teacher-student setting. We have derived exact closed-form equations that describe the training dynamics under online SGD in the high-dimensional limit. By analyzing these equations, we have shown that dropout helps reduce detrimental correlations between the nodes of the network. Furthermore, we have studied the optimal dropout rate and the corresponding generalization error, deriving several analytical results at short, intermediate, and long times. Our findings demonstrate that dropout can significantly mitigate the negative effects of label noise, and that the optimal activation probability decreases as the label noise increases. Finally, we have shown analytically that high dropout rates prevent the network from reaching a specialized configuration.